. A flexible coupling is to be used to connect two shafts, carrying 24 kW power @ 1200 rpm. If the connecting strip is placed 2.5” away from the center of the shafts, how many folds of 0.015” by 0.15” steel made from SAE 1045, quenched in water and drawn at 600oF) are required? (SAE 1045, quenched in water and drawn at 600oF has su = 1 GPa and sy = 775 MPa)

NOTE: Solve this problem (in SI Units) by using a Factor of Safety of 2.5 on su and knowing that the Maximum Shear Stress is the half of the Ultimate Tensile Stress.
Answer #1

First of all let us find the diameter of the shaft that is d.

We know, Mean Torque is given by, T(mean)=60p/2\pi n=60p/2\pi n=24*10^3*60/2\pi *1200=600Nm

where p is the power given and n is the rpm.

Further, there is term called service factor which help us to find tthe maximum torque such that

service factor=1.35

T_m_a_x=1.35*600Nm=810Nm=810*10^3 Nmm

Next , T_m_a_x=\pi /16*\tau *d^3

810*10^3=\pi /16*1*10^6*d^3

d=40mm

D=2d=2*40=80mm

L=1.5d=1.5*40=60mm

Approximately , 28 folds are required.